# SceneVerse: Scaling 3D Vision-Language Learning for Grounded Scene Understanding

### v1

@@ A.2.2 Scene Graph Construction
A.3 Language Generation Details
A.3.1 Object Captioning Pipeline
Object captions aim to provide detailed descriptions of an object’s visual and physical properties, facilitating object-level grounding with its distinctive features. The detailed object captioning pipeline is outlined in Alg. 2. Given the multi-view images $\{I_{1},I_{2},\dots,I_{n}\}$, we utilize the point cloud $P_{o}$ of the object $o$ to get the visible points $P_{o,v}^{vis}$ in the images $v$ through rendering. The occlusion score $s_{o,v}^{occ}$ is calculated as the ratio between the number of visible points and the object point cloud. The image is then cropped with the rendered bounding box and processed through BLIP2 [48] to generate the initial object caption $C_{o,v}$. For each initial caption, we calculate its CLIP [66] similarity score between the text and the cropped image, denoted by $s_{o,v}^{clip}$. To get a refined object caption, we select the top $10$ initial captions with the highest CLIP score and minimal occlusion. The selected sentences are fed into a LLM to obtain a coherent summary of the object captions. In this process, we explicitly instruct the language model to identify and correct the potential errors.
@@ A.3.1 Object Captioning Pipeline
6:         Calculate the similarity score $s^{clip}_{o,v}$ between $C_{o,v}$ and $I^{crop}_{o,v}$ with CLIP [66]
@@
A.3.2 Automatic Language Generation
Template-based
We create diverse templates to generate descriptions for each type of relationship. We categorized the templates into three types based on the number of objects involved and their spatial relationships.
@@ A.3.2 Automatic Language Generation
Description type
Prompt
Object caption
Summarize caption below. The summary should be a description of the target-object. Focus on the target-object’s attribute, like color, shape and material, etc. Identify and correct the potential errors.
@@
Object referral
Rewrite the following caption using one random sentence structure. You should give me only one rewritten sentence without explanation.
caption: The bed is between desk and nightstand.
Rewrite the following caption. You should give me only one rewritten sentence about target-object without explanation. Make sure target-object is the subject of the sentence, not anchor-object(s). If the sentence is in full inversion, keep the inversion.
caption: The armchair is next to the sofa.
target-object: Armchair
anchor-object(s): Sofa
Rewrite the following caption using one random sentence structure. You need to focus on the location and relations of the target-object that appears in the sentence. If multiple target-object appear in the sentence, you need to focus on the first target-object that appears. You can also add the target-object’s function and comfort level based on the sentence, e.g., how the objects can be used by humans and human activities in the scene. You should give me only one rewritten sentence without explanation.
caption: Far from the bowl and peppershaker, the vase is to the left, it is also on the top of countertop.
target-object: Vase
Scene captioning
Your task is to provide a summary for a scene from a given scene graph. The scene contains some objects, which compose a scene graph in json format.
@@
scene graph: { ’scene_type’: ’Bedroom’, ’object_count’: {’nightstand’:2, ...}, ’relation’: {’nightstand’, ’on’, ’floor’}, {’backback’, ’in front of’, bed}, ...}
LLM-rephrasing
To increase description diversity we use the GPT-3.5[61] and Llama[75] for description rephrasing. This improves the diversity and naturalness of the template-based descriptions, as is shown in Fig. 2. The detailed prompts are provided in Tab. A.2.
@@
Appendix B Model Details
B.1 Spatial-Attention Transformer
@@ B.2 Pre-training Details
Appendix C Experimental Details
@@
As briefly mentioned in Sec. 5.1, we mainly considered three model settings in 3D visual grounding experiments, namely scratch, pre-train, and fine-tuned. For the pre-train setting, we follow the same setting mentioned in Sec. B.2. In the scratch and fine-tuned settings, to fairly compare with other dataset-specific fine-tuned models, we add an additional 2-layer MLP over the object features from the referral grounding self-attention transformer. During training, we fine-tune this grounding head together with all model weights for 100 epochs with a learning rate of $1\times 10^{-4}$ for the added projection layer and set all other settings the same as the implementation described in Sec. B.2.
C.2 Zero-shot Transfer
@@
In the zero-shot experiments, we first construct the held-out test set by aggregating scene-text pairs in SceneVerse from scenes in ScanNet and MultiScan. Specifically, we use the validation set of ScanRefer, Nr3D, and Sr3D. For scene-text pairs in the SceneVerse-val, we construct the test set by randomly sampling $\frac{1}{5}$ of human-annotated object referrals in the MultiScan dataset. This results in a test set with around 1.7K object referrals randomly drawn from 8.5k human-annotated object referrals in the MultiScan dataset. In the zero-shot settings, we use all scene-text pairs from datasets in SceneVerse except for ScanNet and MultiScan. This includes both human-annotated and generated texts in ARKitScenes, 3RScan, and HM3D. This setting serves to test models’ generalization capability in grounding objects with both unseen scenes and unseen texts. In the zero-shot text setting, we add generated scene-text pairs in ScanNet and MultiScan into the data used in the zero-shot setting, thereby making the held-out test containing mainly unseen object referrals.
@@
In the zero-shot experiments, we mainly considered three model settings scratch, zero-shot, and zero-shot text. For the zero-shot setting, we pre-train the model following Sec. B.2 without additional grounding heads considering there is no additional training data available in the zero-shot transfer setting. In the scratch and zero-shot text setting, we follow the model implementation described in Sec. C.1 and add an additional 2-layer MLP over the object features from the self-attention transformer. We follow the same fine-tuning setting described in Sec. C.1.
Appendix D Additional Results
In this section, we provide additional experimental results. Specifically, we leverage our collected SceneVerse as the pre-training data source for a traditional 3D semantic segmentation task. Next, we provide ablative analyses of our model design.
D.1 Semantic Segmentation
@@
Table A.3: Semantic segmentation results on ScanNet validation set. normal-$\dagger$ denotes model trained with surface normals as an additional input. S3D indicates models initialized with the original Swin3D model weights pre-trained on Structured3D provided by Yang et al. [85].
@@
D.1.1 Model Ablation
In this section, we provide ablative analyses of our multi-level contrastive alignment design. We mainly consider removing objectives in our model as ablations. We choose the referral-object-level alignment objective as the default setting and consider removing: (i) object-level alignment objective, (ii) masked language modeling objective, and (iii) scene-level alignment objective. For removing the object-level alignment objective, we remove the first stage pre-training of the object point cloud encoder and jointly learn this module within the referral-object-level alignment. As shown in Tab. A.4, we test different model settings on the SceneVerse-val without additional fine-tuning. First, we show that the scene-level alignment objective is crucial for referral object grounding in SceneVerse-val with the $\sim$5performance drop. Similar observations could be made for the model trained without object-level alignment ($\sim$2drop) and masked language modeling objective ($\sim$1.5drop). We believe these results affirm the effectiveness of our overall model design.
Table A.4: Model ablation of our model on SceneVerse-val.
Obj-lvl MLM Scene-lvl Overall Easy Hard
64.8 75.4 48.7
65.2 77.1 47.4
62.4 73.4 45.8
66.9 77.8 50.3

### v3

@@ 0.A.3.1 Object Captioning Pipeline
16:        Calculate the similarity score $s^{clip}_{o,v}$ between $C_{o,v}$ and $I^{crop}_{o,v}$ with
CLIP [78]
@@
0.A.3.2 Automatic Language Generation
  1. 1.
    Template-based We create diverse templates to generate descriptions for each type of relationship. We categorized the templates into three types based on the number of objects involved and their spatial relationships.
@@ 0.A.3.2 Automatic Language Generation
    Description type Prompt
    Object caption Summarize caption below. The summary should be a description of the target-object. Focus on the target-object’s attribute, like color, shape and material, etc. Identify and correct the potential errors.
@@
    Object referral Rewrite the following caption using one random sentence structure. You should give me only one rewritten sentence without explanation. caption: The bed is between desk and nightstand.
    Rewrite the following caption. You should give me only one rewritten sentence about target-object without explanation. Make sure target-object is the subject of the sentence, not anchor-object(s). If the sentence is in full inversion, keep the inversion. caption: The armchair is next to the sofa. target-object: Armchair anchor-object(s): Sofa
    Rewrite the following caption using one random sentence structure. You need to focus on the location and relations of the target-object that appears in the sentence. If multiple target-object appear in the sentence, you need to focus on the first target-object that appears. You can also add the target-object’s function and comfort level based on the sentence, e.g., how the objects can be used by humans and human activities in the scene. You should give me only one rewritten sentence without explanation. caption: Far from the bowl and peppershaker, the vase is to the left, it is also on the top of countertop. target-object: Vase
    Scene captioning Your task is to provide a summary for a scene from a given scene graph. The scene contains some objects, which compose a scene graph in json format.
@@
    scene graph: { ‘scene_type’: ‘Bedroom’, ‘object_count’: {‘nightstand’:2, ...}, ‘relation’: {‘nightstand’, ‘on’, ‘floor’}, {‘backback’, ‘in front of’, ‘bed’}, ...}
  2. 2.
    LLM-rephrasing To increase description diversity we use the GPT-3.5[73] and Llama[86] for description rephrasing. This improves the diversity and naturalness of the template-based descriptions, as is shown in Fig. 2. The detailed prompts are provided in Tab. A.2.
@@
Appendix 0.B Model Details
0.B.1 Spatial-Attention Transformer
@@ 0.B.2 Pre-training Details
Appendix 0.C Experimental Details
@@
As briefly mentioned in Sec. 5.1, we mainly considered three model settings in 3D visual grounding experiments, namely scratch, pre-train, and fine-tuned. For the pre-train setting, we follow the same setting mentioned in Sec. 0.B.2. In the scratch and fine-tuned settings, to fairly compare with other dataset-specific fine-tuned models, we add an additional 2-layer MLP over the object features from the referral grounding self-attention transformer. During training, we fine-tune this grounding head together with all model weights for 100 epochs with a learning rate of $1\times 10^{-4}$ for the added projection layer and set all other settings the same as the implementation described in Sec. 0.B.2.
0.C.2 Zero-shot Transfer
@@
In the zero-shot experiments, we first construct the held-out test set by aggregating scene-text pairs in SceneVerse from scenes in ScanNet and MultiScan. Specifically, we use the validation set of ScanRefer, Nr3D, and Sr3D. For scene-text pairs in the SceneVerse-val, we construct the test set by randomly sampling $\frac{1}{5}$ of human-annotated object referrals in the MultiScan dataset. This results in a test set with around 1.7K object referrals randomly drawn from 8.5k human-annotated object referrals in the MultiScan dataset. In the zero-shot settings, we use all scene-text pairs from datasets in SceneVerse except for ScanNet and MultiScan. This includes both human-annotated and generated texts in ARKitScenes, 3RScan, and HM3D. This setting serves to test models’ generalization capability in grounding objects with both unseen scenes and unseen texts. In the zero-shot text setting, we add generated scene-text pairs in ScanNet and MultiScan into the data used in the zero-shot setting, thereby making the held-out test contain mainly unseen object referrals.
@@
In the zero-shot experiments, we mainly considered three model settings scratch, zero-shot, and zero-shot text. For the zero-shot setting, we pre-train the model following Sec. 0.B.2 without additional grounding heads considering there is no additional training data available in the zero-shot transfer setting. In the scratch and zero-shot text setting, we follow the model implementation described in Sec. 0.C.1 and add an additional 2-layer MLP over the object features from the self-attention transformer. We follow the same fine-tuning setting described in Sec. 0.C.1.
0.C.3 3D question answering
Setting
In the 3D-QA experiments, we evaluate all models with only the training sets provided for fine-tuning. Following previous works [112], we report model performance on the validation and test sets of ScanQA and the test set of SQA3D.
Implementation
As briefly mentioned in Sec. 5.3, we mainly considered the fine-tuned GPS when comparing with existing methods. For all datasets, we initialize our GPS model from a checkpoint pre-trained with 3D visual grounding on SceneVerse. We follow 3D-VisTA and add an additional question-answering module over the pre-trained representations for the answer prediction. We fine-tune the model for 100 epochs with a learning rate of $1\times 10^{-4}$ for the added question answering head and set all other settings the same as the implementation described in Sec. 0.B.2.
0.C.4 Open-vocabulary 3D semantic segmentation
Setting
Following RegionPLC [97] proposed by Yang et al., we conduct experiments to assess the performance of SceneVerse on open-vocabulary 3D semantic segmentation (OV-Seg). To establish a benchmark for open-vocabulary semantic segmentation, we adopt the experimental setup outlined in PLA [31], as per the methodology of RegionPLC. We utilize the annotation-free training setting, as described in RegionPLC, wherein semantic labels for all categories are omitted. This approach allows us to evaluate the effectiveness of SceneVerse in facilitating open-vocabulary segmentation without relying on predefined semantic segmentation annotations. For evaluation, we compute the mean Intersection over Union (mIoU) and mean accuracy (mAcc) across 17 foreground categories, excluding “wall” and “floor” background classes, as well as the “other furniture” category due to its inherent ambiguity.
Implementation
We employ SparseUNet [36] as our 3D backbone network for extracting point features. We utilize different variants of SparseUNet, varying the number of channels in the input layer to explore its impact on performance. For text feature extraction, we employ CLIP [40] text encoder. To align the extracted features from the 3D scene encoder and the text encoder, we incorporate a vision-language adapter. The only supervision comes from the point-discriminative contrastive loss proposed by RegionPLC. During training, we employ the AdamW optimizer to update model parameters. We train the model from scratch for 500 epochs, utilizing a learning rate of $1\times 10^{-3}$. Additionally, we incorporate a warm-up period of 200 steps and a cosine annealing scheme for learning rate scheduling, with a minimum learning rate ratio of $1\times 10^{-5}$.
Appendix 0.D Additional Results
0.D.1 Semantic Segmentation
@@
Table A.3: Semantic segmentation results on ScanNet validation set. $\dagger$ denotes model trained with surface normals as an additional input. S3D indicates models initialized with the original Swin3D model weights pre-trained on Structured3D provided by Yang et al. [99].
@@
0.D.2 Qualitative Results
We provide the qualitative results of 3D vision-language grounding in Fig. A.5 and the results of open-vocabulary semantic segmentation in Fig. A.6.
Figure A.5: Qualitative results of GPS on 3D visual-language grounding. We visualize the incorrect predictions in red and the correct predictions or ground truths in green.
Figure A.6: Qualitative results of open-vocabulary 3D semantic segmentation (OV-Seg). We label the object class in ScanNet-20’s vocabulary in blue, and unseen class in ScanNet-20 in green.